# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ IV-C2 Haptic Communications
The challenges of haptic communications lie in three aspects. The first is the huge communication load caused by the haptic sensors. To guarantee the stability and user experience of teleoperation systems, many types and quantities of haptic sensors should be deployed and they obtain and transmit data at a high rate, e.g., exceeding 1 kHz [273], which brings huge communication overhead. Besides, remote immersive applications, especially critical missions such as telesurgery, are very sensitive to data loss and latency [274]. In summary, haptic communications demand ultra-high reliability and low latency [275].
@@ IV-D Integrated Mobile Crowdsensing and Communication
Typical mobile crowdsensing services are implemented on central servers or clouds. However, the collection of a huge amount of sensory data from massive mobile devices causes a large communication overhead. Besides, most mobile devices communicate via hostile wireless links with channel fading, noise, and limited network resources. Furthermore, many crowdsensing services demand real-time decisions and fast responses, resulting in more severe communication bottlenecks. To maximize the overall crowdsensing system utility calls for communication techniques to overcome the above drawbacks. A direct approach adopts the existing communication schemes in telecommunication networks. They are divided into the edge-network-level techniques and the core-network-level techniques [33]. The former collects sensory data in local wireless edge networks via utilizing existing protocols, including ZigBee, Bluetooth, WiFi, and cellular communications. The latter targets remote data transmission e.g., from mobile devices to clouds, by using a high-speed fiber-optic architecture to provide connections among different wireless edge networks and clouds. Although directly adopting existing communication schemes has low design complexity and is cost-effective, the separated design between communication and crowdsensing degrades the resource utility and cannot fully unleash the performance potential [281]. To this end, a joint design that integrates crowdsensing and communication from a systematic view is desirable. There is currently little work focusing on this area, leaving many research opportunities.
V Signal Design for ISCC
V-A Overview
The above partial-integration techniques, i.e., ISC, ICC, and ISAC, fall short in three aspects as mentioned in Section I-B, i.e., low resource utilization caused by partial integration, mismatch between goals of the overall task and each module caused by separate designs, and ignoring the tight coupling of different modules. To overcome these shortages for improving system performance, ISCC designs are desirable.
@@ V-A Overview
To support ISCC, the signal designs have been investigated including the beamforming and waveform designs. Specifically, the waveforms need to be designed to balance the radar sensing performance evaluated by CRB, the communication performance evaluated by SNR, and the computation performance evaluated by AirComp mean-squared-error (MSE) between the aggregated data and the ground truth. The beamforming designs are further required to equalize the channels among mobile devices and suppress noise to improve the ISCC performance.
@@
V-B Single-Functional Signal Design
@@
By taking the sensing process into consideration, the dedicated signals for sensing, communication, and computation are transmitted simultaneously from multiple single-antenna MDs in [282]. The sensing signals are reflected by the targets and then received by the server, based on which the information of targets is estimated. The performance of sensing is evaluated by the MSE between the estimated target information and the ground truth. The computation signals carrying local learning results at each sensor are aggregated at the server to update the global learning model, where the performance is evaluated by the MSE between the aggregated local model parameters and the ground truth. To reduce the interference between different functionalities, dedicated beamforming designs for sensing, communication, and computation have been investigated. However, the interference among different signals still exists. To reduce the interference, a series of works explore the possibility to integrate the functionalities of sensing, communication, and computation in one common signal.
V-C Dual-Functional Signal Design
In the dual-functional signal design, the antenna array on each MD is divided to produce two distinct signals, one for sensing purposes and the other for AirComp [45]. The information of the sensing target is extracted from the echo signal received at the MD. At the same time, the gathered data is computed through the analog addition of signal waves during the data transmission from the MDs to the server. The MSE of the estimated target response matrix is used for evaluating the sensing performance, while the MSE of the computation results is used for evaluating the AirComp performance. To improve the AirComp accuracy while guaranteeing the sensing performance, the sensing transmit beamformers, data transmit beamformers, and data aggregation beamformers are jointly designed. The impact of AirComp signal interference on radar sensing can be mitigated using statistical methods, while the interference of sensing signals on AirComp will result in deteriorated performance, which needs to be further dealt with.
V-D Triple-Functional Signal Design
@@
VI Network Resource Management for ISCC
VI-A Overview
@@
VI-B Joint RRM for Task Coexistence
@@
VI-B1 Coexistence of Sensing, Communication, and Computation Tasks
@@
In this case, sensing, communication, and computation tasks coexist and compete for network resources to achieve their own targets. Particularly, in [288], ISAC serves as a supporting technique to allow hardware and spectrum sharing between the communication and the sensing tasks. A Cobb-Douglas utility function that combines three metrics, i.e., sensing precision, communication capacity, and computation capability, is adopted and maximized by the joint design of device association and sub-channel assignment. Besides, the authors in [289] investigate a time-division multiplexing ISCC system, where FMCW is used as the sensing waveform and orthogonal time slots are assigned to the sensing, uplink/downlink communication, and computation processes. Targeting communication capacity maximization and computation offloading latency minimization while guaranteeing the sensing distortion is under a limited threshold, a joint resource allocation, time partitioning scheme, computation task processing mode selection, and target sensing location selection, are proposed. In [290], the IRS is applied to assist the ISAC and computation task offloading, where both partial offloading and binary offloading are considered. A performance metric called weighted throughput capacity, which comprises the radar sensing capacity, communication throughput, and computation throughput, is proposed and maximized via joint beamforming design and time allocation. Furthermore, in [28], the ISCC framework is designed for intelligent machine-type communication (IMTC) networks to achieve highly reliable communication with low latency, highly accurate sensing, and fast data processing. In IMTC, intelligent machines (IMs) are expected to execute complex data processing tasks in rapidly varying environments, which demands IMs to conduct real-time environmental sensing and cooperation with each other [291]. To this end, a closed-loop is formulated where the sensed data is exploited as a priori information to improve the communication performance, the communication techniques are designed to support the intelligent computation, and the computation results in turn facilitate the ISAC process.
VI-B2 Coexistence of Sensing and MEC Tasks
@@
In this case, the sensing and the MEC tasks compete for radio resources to achieve their respective goals. ISAC technique is used to enable hardware and radio resource sharing between sensing and communication. In [88], a UAV-supported ISCC framework is proposed, where a treble-functional UAV senses a target by using the unified beam waves, processes sensory data locally, and offloads its computational tasks to ground APs. By characterizing the Pareto boundary between the sensing beampattern gain and the computation capacity, a fundamental trade-off is revealed. Specifically, by applying the semidefinite programming method and the concave-convex procedure technique, a joint optimization algorithm is proposed and depicts the performance region, as well as enhancing the ability of sensing, computing, and communication of UAVs. Besides, an architecture that integrates communication, radar sensing, and MEC, is developed in [292], where user terminals conduct computation offloading and radar sensing over the same radio resource blocks by using DFRC techniques and MIMO arrays. To jointly optimize the performance of MIMO radar beampattern design, computation offloading energy consumption, the individual transmit precoding for radar, and resource allocation of communication and computation, a multi-objective optimization problem is formulated. Moreover, the authors in [293] develop a non-orthogonal multiple access (NOMA)-aided joint communication, sensing, and multi-tier computing (JCSMC) framework, where the BS can not only sense target areas and offer MEC services to the covered mobile devices simultaneously but also offload some computation tasks to cloud servers to improve computation efficiency. The potential benefits of utilizing NOMA in this framework are investigated. Both the transmit beamformer and computing resource allocation are jointly optimized to maximize the computation rate in two computation offloading modes, i.e., partial and binary offloading modes.
VI-B3 Coexistence of ISC and MEC Tasks

### v2

@@ IV-C2 Haptic Communications
The challenges of haptic communications lie in three aspects. The first is the huge communication load caused by the haptic sensors. To guarantee the stability and user experience of teleoperation systems, many types and quantities of haptic sensors should be deployed and they obtain and transmit data at a high sampling rate, e.g., exceeding 1 kHz [264], which brings huge communication overhead. Besides, remote immersive applications, especially critical missions such as telesurgery, are very sensitive to data loss and latency [265]. In summary, haptic communications demand ultra-high reliability and low latency [266].
@@ IV-D Integrated Mobile Crowdsensing and Communication
Typical mobile crowdsensing services are implemented on central servers or clouds. However, the collection of a huge amount of sensory data from massive mobile devices causes a large communication overhead. Besides, most mobile devices communicate via hostile wireless links with channel fading, noise, and limited network resources. Furthermore, many crowdsensing services demand real-time decisions and fast responses, resulting in more severe communication bottlenecks. To maximize the overall crowdsensing system utility calls for communication techniques to overcome the above drawbacks. A direct approach adopts the existing communication schemes in telecommunication networks. They are divided into the edge-network-level techniques and the core-network-level techniques [34]. The former collects sensory data in local wireless edge networks via utilizing existing protocols, including ZigBee, Bluetooth, WiFi, and cellular communications. The latter targets remote data transmission e.g., from mobile devices to clouds, by using a high-speed fiber-optic architecture to provide connections among different wireless edge networks and clouds. Although directly adopting existing communication schemes has low design complexity and is cost-effective, the separated design between communication and crowdsensing degrades the resource utility and cannot fully unleash the performance potential [272]. To this end, a joint design that integrates crowdsensing and communication from a systematic view is desirable. There is currently little work focusing on this area, leaving many research opportunities. One possible research direction lies in adaptively designing the data sensing, collection, and processing approaches according to the time-varying wireless channels and communication capacities. Another possible research direction is joint resource allocation among the processes of sensing, communication, data filtering, and quantization to maximize the data utility.
IV-E Discussion
The three ISAC paradigms have their own advantages and challenges. The paradigm of integrated wireless sensing and communication enjoys the benefit of low cost since it can reuse the ubiquitous signals and hardware in wireless communication networks for sensing. Besides, the techniques of wireless signal processing and transceiver design are mature. Moreover, the existing methods in communication systems, such as MIMO beamforming and OFDM modulation, can be directly implemented in this paradigm. These make it easy to implement. However, the interference management between sensing and communication as well as the requirements to achieve both goals of the two tasks bring new design challenges. In the paradigm of integrated multi-modal sensing and communication, communication enables cooperative multi-modal sensing among many sensors to complete complicated tasks and achieve higher task performance. On the other hand, multi-modal sensing can provide abundant prior knowledge for enhancing communication quality. The challenges lie in two aspects. One is that the multi-modal sensory data are in various formats and differ in physical properties like resolution, bandwidth, and latency. The other is the lack of alignment, labeling, and fusion techniques for constructing practical datasets. For integrated mobile crowdsensing and communication, it enjoys the benefit of collecting massive data for completing many tasks to improve the data utility. The challenges arise from the data privacy, the huge communication load, and the characterization of the data utility function.
V Benefits, Functions, and Challenges of ISCC
ISCC refers to the integrated design of sensing, communication, and computation modules via resources (e.g., energy, time, spectrum, and hardware) sharing and coordination, and joint algorithm development among them. In thi section, the benefits of ISCC, the functions of ISCC in 6G, and its challenges are discussed.
V-A Benefits
Compared with existing partial integration technologies, ISCC enjoys the benefits of sensing-communication-computation symbiosis and overcoming the three shortages mentioned in Section I-B to achieve higher resource utilization and task performance.
Figure 7: Sensing-Communication-Computation Symbiosis in ISCC systems.
To begin with, ISCC has the opportunity for sensing, communication, and computation symbiosis, as shown in Fig. 7. Specifically, the information obtained via sensing can be used as prior knowledge to enhance communication quality and the data resources for executing complicated computational tasks, e.g., federated learning and edge AI inference. Communication serves as an information-sharing pipeline for cooperative sensing and distributed computational tasks. Computation can provide intelligent and efficient execution of algorithms for enhancing the sensing and communication performance. This ISCC-enabled closed-loop framework can achieve mutual performance improvement in all three modules.
Then, ISCC can achieve higher resource utilization than the three partial integration technologies. In the networks that are enabled with partial integration technologies, network resources are partially shared between two modules or even orthogonally shared among the three modules. For example, in ISAC systems, radio and hardware resource sharing is enabled between communication and sensing tasks, but not for computation tasks [36]. However, in an ISCC system, the waveforms corresponding to the sensing communication, and computation tasks can coexist in the same radio resource block [273] or one waveform is directly designed for completing the three kinds of tasks [274]. On the other hand, compared with existing partial integration technologies, where network resources are partially coordinated between one or two modules, network resource management among sensing, communication, and computation modules from a systematic view is enabled. For example, in ISCC-enabled edge-device cooperative systems, ISCC resource allocation can save resource costs to achieve the same performance compared with separate resource allocation schemes [12, 42].
Figure 8: The sensing, computation, and communication processes in complicated tasks.
Moreover, for complicated tasks that involve sensing, computation, and communication processes as shown in Fig. 8, existing partial integration technologies face shortages of mismatch between the goals of the overall task and each module and ignoring the tight coupling of different modules, as mentioned in Section I-B. The former shortage refers to the design metrics of each module in existing partial integration technologies, e.g., mutual information for sensing, throughput for communication, and MSE for AirComp, don’t match the goal of the overall complicated task. As an example, in an edge AI inference task, enhancing the sensing quality of the least important view or transmitting the irrelevant feature elements may enhance the traditional design goals of sensing and communication but is useless for improving the inference accuracy. The latter shortage refers to existing partial integration technologies failing to respect the coupling mechanism where the task performance depends on the quality of each module but different modules compete for network resources. For example, in an ISCC-supported online federated learning framework proposed in [275], the convergence rate is decided by the number of sensory data samples, the sensing SNR, the communication SNR, and the computation speed, but the three modules demand to compete for time and energy to improve their own higher qualities. ISCC can overcome the above two shortages to achieve higher task performance by adaptively sharing, coordinating, and allocating network resources among sensing, communication, and computation in a task-oriented way, i.e., under a unified task goal.
V-B ISCC’s Functions in 6G
ISCC will be a key enabler to enhance system performance in 6G and beyond networks. In the sequel, we shall discuss how ISCC can enhance 6G KPIs and facilitate the usage scenarios.
V-B1 ISCC Enhances 6G KPIs
With the advantages of sensing-communication-computation symbiosis, higher resource utilization, and better task performance, ISCC can help achieve many 6G KPIs. Some examples are listed below.
  • Spectrum Efficiency: ISCC enables different kinds of waveforms (i.e., sensing, communication, and computation) to share the same frequency band [276] or directly utilize one waveform to complete the three functions [277]. Besides, in ISCC systems, AI methods can be utilized to process real-time sensory data for spectrum detection and prediction [278]. As a result, the spectrum efficiency is improved.
  • Mobility: By carefully designing the multi-modal sensing strategy toward moving communication devices, their velocity, positions, and trajectories can be estimated or predicted. This can help compensate for the Doppler effect and enable soft handoff among different cells, especially for fast-moving devices.
  • Task-completion Latency: By adaptively allocating the time among sensing, communication, and computation modules and allowing pipeline design for long-run tasks, e.g., allowing time overlapping between sensing for obtaining a new data sample and the extraction and the transmission of the local feature vectors in continuous AI inference tasks, the end-to-end task completion latency is saved.
  • Connection Density: In a dense network, the devices’ locations and channel gains can be estimated by intelligently processing the real-time multi-modal sensory data. Therefore, better beamforming, time, or subcarrier allocation strategies can be designed to enhance the connectivity.
  • Energy Efficiency: ISCC can help enhance the network energy efficiency. On one hand, it allows reusing signals from one transmitter for completing multiple tasks of sensing, communication, and computation. On the other hand, the energy of devices and servers can be adaptively managed among the three processes.
V-B2 ISCC for 6G Usage Scenarios
As a key technology in 6G, ISCC is expected to support a series of usage scenarios, including edge intelligence, metaverse, IoT networks, smart city, smart ocean, satellite, machine system, SAGIN, vehicular networks, and Wi-Fi sensing [279]. As for edge intelligence, ISCC enables simultaneous high-quality environmental data sensing, delivery, and computation to maximize the inference/training accuracy under constraints on low latency and on-device resources [12]. The task-oriented ISCC designs are further tailored to account for different AI applications [280]. As an AI-generated virtual world, the metaverse requires tremendous sensing, communication, and computation resources, where the ISCC can be applied to cope with the conflict between limited resources and user demands [50]. In IoT networks, the application of ISCC can improve the spectrum efficiency and reduce latency by enabling simultaneous sensing, communication, and computation [281]. In smart city, ISCC is expected to promote the development of information infrastructure [282]. In smart ocean, ISCC is applied for connecting marine objects in surface and underwater environments [283]. As for satellites, ISCC is proposed to cope with the complex and dynamic satellite environments and limited network resources [284]. In machine systems, the application of ISCC can realize low latency and high reliability of communication, highly accurate sensing, and fast environment adaption [29]. In SAGIN, ISCC is applied to meet the quality of service requirements of various applications [285]. As for vehicular networks, ISCC is applied to increasing the sensing accuracy of the vehicular-equipped sensors while reducing the data communication and computation latency [286, 287, 288]. ISCC is also expected to be deployed in the scenario of Wi-Fi sensing for improving the data processing efficiency [289].
V-C Challenges
Although the ISCC design can achieve a higher performance ceiling, it encounters new challenges, as elaborated in the following.
  • Multi-objective Optimization: In the case of multiple-task coexistence, different tasks have different types of design targets and metrics. However, it is challenging to design a way that can jointly optimize these design metrics since they have different units. For example, the communication throughput is measured by bits, AirComp performance is generally measured by MSE between the computation result and ground truth, and sensing is measured by CRB. Existing designs generally combine these design metrics in a weighted sum form with a manually pre-determined weight for each metric. This fails to accurately characterize the importance levels of different tasks, leading to a sub-optimal performance. It remains as a challenge to design an efficient and effective multi-objective optimization mechanism.
  • Generalization: Existing ISCC schemes are designed for specific tasks (e.g. federated learning, edge AI inference), for certain networks (vehicular or sensor networks), or under particular scenarios (coexistence of two or more specific tasks). There is a lack of an ISCC strategy that can be commonly utilized or easily transferred to other tasks, networks, or scenarios.
Many efforts have been put on ISCC to address these challenges. In this paper, we categorize these techniques into two types, i.e., signal design and network resource management, which are presented in Sections VI and VII respectively.
VI Signal Design for ISCC
VI-A Overview
@@
To support ISCC, the signal designs have been investigated including the beamforming and waveform designs. Specifically, the waveforms need to be designed to balance the radar sensing performance evaluated by CRB, the communication performance evaluated by SNR, and the computation performance evaluated by AirComp MSE between the aggregated data and the ground truth. The beamforming designs are further required to equalize the channels among mobile devices and suppress noise to improve the ISCC performance.
@@
VI-B Multi-Functional Signal Design
VI-B1 Single-Functional Signal Design
@@
By taking the sensing process into consideration, the dedicated signals for sensing, communication, and computation are transmitted simultaneously from multiple single-antenna MDs in [276]. The sensing signals are reflected by the targets and then received by the server, based on which the information of targets is estimated. The performance of sensing is evaluated by the MSE between the estimated target information and the ground truth. To reduce the interference between different functionalities, dedicated beamforming designs for sensing, communication, and computation have been investigated in [276]. The algorithm complexity of the beamforming design is $\sqrt{2K}(n_{1}K+n_{1}^{3})$, where $n_{1}=\mathcal{O}(K^{2}M^{2})$ with $K$ denoting the number of MDs and $M$ denoting the number of antennas at each MD. The computation signals carrying local learning results at each sensor are aggregated at the server to update the global learning model, where the performance is evaluated by the MSE between the aggregated local model parameters and the ground truth. To reduce the interference between different functionalities, dedicated beamforming designs for sensing, communication, and computation have been investigated. However, the interference among different signals still exists. To reduce the interference, a series of works explore the possibility of integrating the functionalities of sensing, communication, and computation in one common signal.
VI-B2 Dual-Functional Signal Design
In the dual-functional signal design, the antenna array on each MD is divided to produce two distinct signals, one for sensing purposes and the other for AirComp [274]. The information of the sensing target is extracted from the echo signal received at the MD. At the same time, the gathered data is computed through the analog addition of signal waves during the data transmission from the MDs to the server. The MSE of the estimated target response matrix is used for evaluating the sensing performance, while the MSE of the computation results is used for evaluating the AirComp performance. Compared with the single-functional design where the interference between the signals for sensing, communication and computation needs to be managed, the dual-functional design only needs to manage the interference between the signals for sensing and AirComp. To improve the AirComp accuracy while guaranteeing the sensing performance, the sensing transmit beamformers, data transmit beamformers, and data aggregation beamformers are jointly designed. The impact of AirComp signal interference on radar sensing can be mitigated using statistical methods, while the interference of sensing signals on AirComp will result in deteriorated performance, which needs to be further dealt with.
VI-B3 Triple-Functional Signal Design
@@
VI-C Discussion
The ISCC signals can be applied in a series of scenarios including target location estimation, vehicular networks, and federated learning [281]. As for target location estimation, the location of the target is estimated by multiple MDs based on the information of relative distance and angle estimated from the reflected radar signals, as well as their own locations. Meanwhile, each MD delivers its previous locally estimated location of the target to the server simultaneously, and thus the server obtains the averaged estimated target location via AirComp. In vehicular networks, the ISCC signals are expected to train the beam for vehicular tracking and predicting. In federated learning, the ISCC signals are used to simultaneously sense the information of the target for local model training and deliver the local training results to the server via AirComp for global model update [280]. The single-functional design has the advantage of simple waveform design for dedicated functionality, but suffers from severe interference. The dual-functional design mitigates the interference between communication and computation interference by exploiting the AirComp property, while the interference between sensing and AirComp still exists. The triple-functional design gets rid of the interference among different functionalities, but at sacrifice of complex waveform design.
VII Network Resource Management for ISCC
VII-A Overview
@@
VII-B Joint RRM for Task Coexistence
@@
VII-B1 Coexistence of Sensing, Communication, and Computation Tasks
@@
In this case, sensing, communication, and computation tasks coexist and compete for network resources to achieve their own targets. Particularly, in [296], ISAC serves as a supporting technique to allow hardware and spectrum sharing between the communication and the sensing tasks. A Cobb-Douglas utility function that combines three metrics, i.e., sensing precision, communication capacity, and computation capability, is adopted and maximized by the joint design of device association and sub-channel assignment. Besides, the authors in [297] investigate a time-division multiplexing ISCC system, where FMCW is used as the sensing waveform and orthogonal time slots are assigned to the sensing, uplink/downlink communication, and computation processes. Targeting communication capacity maximization and computation offloading latency minimization while guaranteeing the sensing distortion is under a limited threshold, a joint resource allocation, time partitioning scheme, computation task processing mode selection, and target sensing location selection, are proposed. In [298], the IRS is applied to assist the ISAC and computation task offloading, where both partial offloading and binary offloading are considered. A performance metric called weighted throughput capacity, which comprises the radar sensing capacity, communication throughput, and computation throughput, is proposed and maximized via joint beamforming design and time allocation. Furthermore, in [29], the ISCC framework is designed for intelligent machine-type communication (IMTC) networks to achieve highly reliable communication with low latency, highly accurate sensing, and fast data processing. In IMTC, intelligent machines (IMs) are expected to execute complex data processing tasks in rapidly varying environments, which demands IMs to conduct real-time environmental sensing and cooperation with each other [299]. To this end, a closed-loop is formulated where the sensed data is exploited as a priori information to improve the communication performance, the communication techniques are designed to support the intelligent computation, and the computation results in turn facilitate the ISAC process. Moreover, the resource allocation among sensing, communication, and computation tasks in vehicular networks is investigated in [300] under the principle of maximizing the achievable data rate while guaranteeing sensing and computation performance.
VII-B2 Coexistence of Sensing and MEC Tasks
@@
In this case, the sensing and the MEC tasks compete for radio resources to achieve their respective goals. ISAC technique is used to enable hardware and radio resource sharing between sensing and communication. In [301], a UAV-supported ISCC framework is proposed, where a treble-functional UAV senses a target by using the unified beam waves, processes sensory data locally, and offloads its computational tasks to ground APs. By characterizing the Pareto boundary between the sensing beampattern gain and the computation capacity, a fundamental trade-off is revealed. Specifically, by applying the semidefinite programming method and the concave-convex procedure technique, a joint optimization algorithm is proposed and depicts the performance region, as well as enhancing the ability of sensing, computing, and communication of UAVs. Besides, an architecture that integrates communication, radar sensing, and MEC, is developed in [302], where user terminals conduct computation offloading and radar sensing over the same radio resource blocks by using DFRC techniques and MIMO arrays. To jointly optimize the performance of MIMO radar beampattern design, computation offloading energy consumption, the individual transmit precoding for radar, and resource allocation of communication and computation, a multi-objective optimization problem is formulated. The authors in [303] develop a non-orthogonal multiple access (NOMA)-aided joint communication, sensing, and multi-tier computing (JCSMC) framework, where the BS can not only sense target areas and offer MEC services to the covered mobile devices simultaneously but also offload some computation tasks to cloud servers to improve the computation efficiency. The potential benefits of utilizing NOMA in this framework are investigated. Both the transmit beamformer and computing resource allocation are jointly optimized to maximize the computation rate in two computation offloading modes, i.e., partial and binary offloading modes. An IRS-assisted ISCC system is proposed in [304] to support multiple devices simultaneously sensing for target detection and computation offloading. The sensing quality and the communication throughput for computation offloading are enhanced by the IRS. Moreover, a multi-point aided ISCC framework is proposed in [305], where multiple base stations cooperatively conduct target detection via sensing and serve computation offloading for multiple devices simultaneously.
VII-B3 Coexistence of ISC and MEC Tasks